# Eigenpruning

### v1

@@ Abstract
We introduce eigenpruning, a method that removes singular values from weight matrices in an LLM to improve its performance in a particular task. This method is inspired by interpretability methods designed to automatically find subnetworks of a model which solve a specific task. In our tests, the pruned model outperforms the original model by a large margin, while only requiring minimal computation to prune the weight matrices. In the case of a small synthetic task in integer multiplication, the Phi-2 model can improve its accuracy in the test set from 13.75% to 97.50%. Interestingly, these results seem to indicate the existence of a computation path that can solve the task very effectively, but it was not being used by the original model. Finally, we plan to open-source our implementation in the non-anonymous version of our work.
@@ 4 Experiments
We used 2 open-source LLMs: GPT2-XL Radford et al. (2019), Phi-2 Javaheripi et al. (2023). For GPT2-XL, we also finetune it to each task to check whether there is an effect of combining finetuning with eigenpruning. Details can be found in Appendix C.
@@ 7 Limitations
This is work in progress, and many limitations are present. First, we have not tried sufficient models to be able to understand if the potential of eigenpruning is truly generalizable to all LLMs. For example, GPT-2 had much less of a benefit from eigenpruning than Phi-2. It might be the case that eigenpruning just needs more capable models to be useful, but we do not currently have the data to make that claim.
Second, we have not enough data to understand the effect of finetuning with eigenpruning. We only tested finetuning GPT-2, but we need to test more models and finetune each one to actually understand its effects.

### v2

@@ Abstract
We introduce eigenpruning, a method that removes singular values from weight matrices in an LLM to improve its performance in a particular task. This method is inspired by interpretability methods designed to automatically find subnetworks of a model which solve a specific task. In our tests, the pruned model outperforms the original model by a large margin, while only requiring minimal computation to prune the weight matrices. In the case of a small synthetic task in integer multiplication, the Phi-2 model can improve its accuracy in the test set from 13.75% to 97.50%. Interestingly, these results seem to indicate the existence of a computation path that can solve the task very effectively, but it was not being used by the original model. Finally, we publicly release our implementation [^1] https://github.com/tvergara/eigenpruning.
@@ 2 Related Work
  • Other work has shown that using attribution methods to prune attention heads can improve the accuracy in certain tasks Hao et al. (2021).
@@ 4 Experiments
We used 2 open-source LLMs: GPT2-XL Radford et al. (2019), Phi-2 Javaheripi et al. (2023). For GPT2-XL, we also finetune it to each task before applying our method to check whether there is an effect of combining finetuning with eigenpruning. Details can be found in Appendix C.
@@ 7 Limitations
This is work in progress, and many limitations are present.
First, we have not tried sufficient models to be able to understand if the potential of eigenpruning is truly generalizable to all LLMs. For example, GPT-2 had much less of a benefit from eigenpruning than Phi-2. It might be the case that eigenpruning just needs more capable models to be useful, but we do not currently have the data to make that claim.
Second, the optimal $p$ value for each combination of task/model is currently done as a brute force search over several candidate values, which is far from being ideal.
Third, we have not enough data to understand the effect of finetuning with eigenpruning. We only tested finetuning GPT-2, but we need to test more models and finetune each one to actually understand its effects.
Fourth, our work needs comparison to other parameter efficient finetuning methods, such as LoRA Hu et al. (2021), and potentially also with model editing approaches, such as Mitchell et al. (2021).
@@
8 Acknowledgements
We appreciate the reviewers’ feedback on this version of our work. We received suggestions for related work that we were previously not aware about. We believe these might be very influential in our future work in eigenpruning.
Also, we would like to thank LatinX for the opportunity to present our work. They are doing amazing work into providing opportunities for Latinx individuals along the world.
Finally, we would like to thank Aizawa-lab students for their feedback in previous versions of our diagrams, and overall suggestions over the course of this research.